If f( x ) = 5x+3 and g ( x ) = x^2-3 for what val;ue of x is f( x )= g( x )?

Answer by venugopalramana(3286)


If f( x ) = 5x+3 and g ( x ) = x^2-3 for what val;ue of x is f( x )= g( x )?
we have 5x+3=x^2-3
x^2-3-5x-3=0...or ....x^2-5x-6=0
x^2-6x+x-6=0
x(x-6)+1(x-6)=0
(x-6)(x+1)=0
so x-6=0....or.....x+1=0
so x=6.....or x=-1...are the 2 values of x for f( x )= g( x )
--------------------------------------------------------------------------------